# GaussianArt: Unified Modeling of Geometry and Motion for Articulated Objects

### v1

@@ 10 Multi-view Mask Consistency
The fine-tuned model can generate masks at the desired granularity, but the masks are inconsistent across views. Some studies [78, 3] adopt a contrastive learning strategy to achieve multi-view consistency. A random feature vector is initialized for each Gaussian primitive and optimized using a contrastive loss. At test time, object masks are obtained through density-based clustering [8] and projected to the target view. This approach is effective for prompt-based segmentation; however, it is not well suited for part-level motion supervision.
In the context of single-object 3D segmentation, monocular masks provide only partial coverage of the target object, making them insufficient for comprehensive observation. To address this limitation, SA3D [2] introduced a multi-view mask re-prompting algorithm to mitigate the issue of incomplete observations. Building upon this approach, we extend the mask re-prompting method proposed in SA3D [2] to a multi-mask setting to generate multi-view consistent part labels. Given source view and target view , we have RGB-D images , camera parameters and segmentation masks generated by Art-SAM: , where and denote the number of segmentation masks. For each mask obtained from the source view mask, we first obtain a 3D point mask by projecting each of the masked regions with the depth map and camera parameters:
@@
To ensure cross-state consistency, we first obtain 2D pixel correspondences between images at two states using LoFTR [52]. These matches allow us to map segmentation categories between states, aligning part segmentations across them.
@@ 13 More Discussion of Gaussian-to-Mesh
For objects with severely uneven view distributions, such as the USB and Stapler in PARIS [33], the mesh method employed in this work fails to complete regions with sparse views, leading to a low CD. Additionally, objects with noisy depth maps, such as real objects, may result in holes in the extracted mesh. While a tetrahedral grid-based method in [80] can mitigate these issues, it often introduces surface noise when using vanilla 3DGS. In order to optimize surface reconstruction, we attempted to adapt the method from GaussianArt to Gaussian Opacity Fields [80] or impose surface constraints on the vanilla baseline. However, none of these approaches could effectively learn the articulated motion. This, to some extent, indicates that the modeling methods for flattened Gaussians in articulated objects require further exploration. In future research, we will explore more effective mesh reconstruction methods while maintaining accurate motion estimation to improve mesh quality under different conditions.
@@ 15.1 Datasets
PARIS Two-Part Dataset. The dataset created by [33] contains multi-view posed renderings of objects spanning 10 categories in PM, along with 2 real world scans of articulated objects of the same fashion. We follow [65] to use an enhanced version with rendered depth maps.
DigitalTwinArt-PM Dataset. A multi-part dataset proposed by [65], containing two 3-part articulated objects from PM, each with one static part and two movable parts.
@@ 16 Results on Basic Datasets
[…]
| PARIS [33] | 13.2 | 28.3 | 0.0 | 0.0 | 3.4 | 35.1 | 0.2 | 30.3 | 14.9 | 0.0 | 12.5 | 0.3 | 23.4 | 11.9 | |
| PARIS* [33] | 29.3 | 5.3 | 0.0 | 23.4 | 32.1 | 15.3 | 0.8 | 23.3 | 31.4 | 0.0 | 16.1 | 0.0 | 10.4 | 5.2 | |
| CSG-reg [65] | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.0 | 10.5 | 6.3 | 5.0 | 2.5 | 2.4 | 0.9 | 9.6 | 5.2 | |
| 3Dseg-reg [65] | - | - | 0.11 | - | - | - | - | - | 7.5 | - | - | - | - | - | |
[…]
| PARIS [33] | 0.9 | 1.0 | 0.0 | 0.0 | 1.4 | 3.4 | 3.4 | 1.3 | - | - | 1.1 | 0.2 | - | 0.2 | |
| PARIS* [33] | 0.5 | 2.6 | 0.0 | 0.2 | 1.7 | 3.9 | 3.1 | 3.1 | - | - | 1.5 | 0.0 | - | 0.0 | |
| CSG-reg [65] | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 0.0 | 0.0 | 1.5 | - | - | 0.2 | 1.1 | - | 1.1 | |
| 3Dseg-reg [65] | - | - | 0.0 | - | - | - | - | - | - | - | - | - | - | - | |
[…]
| PARIS [33] | 78.9 | 57.7 | 0.0 | 0.0 | 50.1 | 47.1 | 84.3 | 23.3 | 0.2 | 0.0 | 34.2 | 1.1 | 0.4 | 0.7 | |
| PARIS* [33] | 75.0 | 58.5 | 0.0 | 28.8 | 64.8 | 68.1 | 67.8 | 26.5 | 0.2 | 0.0 | 39.0 | 0.0 | 0.1 | 0.1 | |
| CSG-reg [65] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 9.3 | 8.8 | 0.0 | 0.0 | 1.8 | 0.1 | 0.1 | 0.1 | |
| 3Dseg-reg [65] | - | - | 0.1 | - | - | - | - | - | 0.0 | - | - | - | - | - | |
[…]
| PARIS [33] | 5.0 | 2.7 | 0.0 | 1.0 | 3.8 | 0.2 | 0.3 | 2.2 | 0.6 | 2.5 | 1.8 | 34.1 | 30.1 | 32.1 | |
| PARIS* [33] | 5.8 | 12.0 | 0.0 | 0.3 | 13.3 | 1.9 | 0.5 | 3.7 | 0.7 | 0.8 | 3.9 | 1.5 | 2.8 | 2.1 | |
| CSG-reg [65] | |||||||||||||||
| 3Dseg-reg [65] | - | - | - | - | - | - | - | - | - | - | - | - | |||
[…]
| PARIS [33] | 7.6 | 11.2 | 0.2 | 15.2 | 40.7 | 30.7 | 5.9 | 131.7 | 9.5 | 6.1 | 25.9 | 31.7 | 382.9 | 207.3 | |
| PARIS* [33] | 24.9 | 11.5 | 0.0 | 10.6 | 16.6 | 37.0 | 6.9 | 248.7 | 213.0 | 11.0 | 58.0 | 26.8 | 227.2 | 127.0 | |
| CSG-reg [65] | |||||||||||||||
| 3Dseg-reg [65] | - | - | - | - | - | - | - | - | - | - | - | - | |||
[…]
| PARIS [33] | 1.2 | 0.9 | 0.1 | 1.1 | 20.7 | 4.2 | 0.2 | 3.3 | 0.2 | 0.4 | 3.2 | 15.5 | 22.2 | 18.9 | |
| PARIS* [33] | 6.4 | 4.7 | 0.0 | 0.3 | 3.6 | 5.9 | 0.2 | 2.8 | 0.1 | 0.4 | 2.4 | 1.4 | 8.2 | 4.8 | |
| CSG-reg [65] | |||||||||||||||
| 3Dseg-reg [65] | - | - | - | - | - | - | - | - | - | - | - | - | |||
[…]
@@
| PSNR | PARIS [33] | 33.11 | 38.78 | 38.66 | 35.41 | 38.70 | 38.59 | 38.24 | 40.18 | 38.56 | 37.08 | 37.62 | 25.29 | 27.13 | 26.21 |
| DigitalTwinArt [65] | 42.79 | 32.99 | 37.72 | 34.73 | 40.41 | 38.28 | 38.78 | 38.71 | 43.16 | 35.38 | 38.30 | 24.22 | 23.11 | 23.67 | |
| ArtGS [38] | 34.46 | 37.11 | 34.09 | 37.06 | 38.29 | 39.13 | 39.64 | 38.50 | 41.16 | 37.24 | 37.67 | 27.05 | 25.38 | 26.22 | |
@@
| SSIM | PARIS [33] | 0.986 | 0.994 | 0.990 | 0.981 | 0.996 | 0.995 | 0.992 | 0.991 | 0.996 | 0.994 | 0.992 | 0.898 | 0.953 | 0.926 |
| DigitalTwinArt [65] | 0.986 | 0.986 | 0.989 | 0.977 | 0.994 | 0.991 | 0.991 | 0.992 | 0.997 | 0.968 | 0.987 | 0.890 | 0.923 | 0.907 | |
| ArtGS [38] | 0.997 | 0.993 | 0.988 | 0.995 | 0.998 | 0.999 | 0.998 | 0.995 | 0.999 | 0.992 | 0.995 | 0.939 | 0.930 | 0.935 | |
@@
[…]

### v2

@@ 10 Multi-view Mask Consistency
The fine-tuned model can generate masks at the desired granularity, but the masks are inconsistent across views. Some studies [Ying_2024_Omniseg3d, cen2024segment3dgaussians] adopt a contrastive learning strategy to achieve multi-view consistency. A random feature vector is initialized for each Gaussian primitive and optimized using a contrastive loss. At test time, object masks are obtained through density-based clustering [ester1996density] and projected to the target view. This approach is effective for prompt-based segmentation; however, it is not well suited for part-level motion supervision.
In the context of single-object 3D segmentation, monocular masks provide only partial coverage of the target object, making them insufficient for comprehensive observation. To address this limitation, SA3D [cen2023_sa3d] introduced a multi-view mask re-prompting algorithm to mitigate the issue of incomplete observations. Building upon this approach, we extend the mask re-prompting method proposed in SA3D [cen2023_sa3d] to a multi-mask setting to generate multi-view consistent part labels. Given source view and target view , we have RGB-D images , camera parameters and segmentation masks generated by Art-SAM: , where and denote the number of segmentation masks. For each mask obtained from the source view mask, we first obtain a 3D point mask by projecting each of the masked regions with the depth map and camera parameters:
@@
To ensure cross-state consistency, we first obtain 2D pixel correspondences between images at two states using LoFTR [sun2021loftr]. These matches allow us to map segmentation categories between states, aligning part segmentations across them.
@@ 12 Mesh Extraction
To extract meshes from Gaussians, we employ the method of rendering median depth as introduced in [keetha2024splatam] and fuse it into a mesh using VDBFusion [vizzo2022vdbfusion]. This entire process can be efficiently accomplished with GauS [ye2024gaustudio] with appropriate selections of voxel size and truncated threshold.
@@ 13 More Discussion of Gaussian-to-Mesh
For objects with severely uneven view distributions, such as the USB and Stapler in PARIS [liu2023paris], the mesh method employed in this work fails to complete regions with sparse views, leading to a low CD. Additionally, objects with noisy depth maps, such as real objects, may result in holes in the extracted mesh. While a tetrahedral grid-based method in [yu2024gaussian] can mitigate these issues, it often introduces surface noise when using vanilla 3DGS. In order to optimize surface reconstruction, we attempted to adapt the method from GaussianArt to Gaussian Opacity Fields [yu2024gaussian] or impose surface constraints on the vanilla baseline. However, none of these approaches could effectively learn the articulated motion. This, to some extent, indicates that the modeling methods for flattened Gaussians in articulated objects require further exploration. In future research, we will explore more effective mesh reconstruction methods while maintaining accurate motion estimation to improve mesh quality under different conditions.
@@ 15.1 Datasets
PARIS Two-Part Dataset. The dataset created by [liu2023paris] contains multi-view posed renderings of objects spanning 10 categories in PM, along with 2 real world scans of articulated objects of the same fashion. We follow [weng2024digitaltwin] to use an enhanced version with rendered depth maps.
DigitalTwinArt-PM Dataset. A multi-part dataset proposed by [weng2024digitaltwin], containing two 3-part articulated objects from PM, each with one static part and two movable parts.
@@ 16 Results on Basic Datasets
| Simulation | Real | ||||||||||||||
| FoldChair | Fridge | Laptop† | Oven† | Scissor | Stapler | USB | Washer | Blade | Storage† | All | Fridge | Storage | All | ||
| Axis Ang | Ditto [Jiang_2022_ditto] | * | * | =20.74pt,depth=17.28pt,width=0.0pt | |||||||||||
| PARIS [liu2023paris] | 13.2 | 28.3 | 0.0 | 0.0 | 3.4 | 35.1 | 0.2 | 30.3 | 14.9 | 0.0 | 12.5 | 0.3 | 23.4 | 11.9 | |
| PARIS* [liu2023paris] | 29.3 | 5.3 | 0.0 | 23.4 | 32.1 | 15.3 | 0.8 | 23.3 | 31.4 | 0.0 | 16.1 | 0.0 | 10.4 | 5.2 | |
| CSG-reg [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.0 | 10.5 | 6.3 | 5.0 | 2.5 | 2.4 | 0.9 | 9.6 | 5.2 | |
| 3Dseg-reg [weng2024digitaltwin] | - | - | 0.11 | - | - | - | - | - | 7.5 | - | - | - | - | - | |
| DigitalTwinArt [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 =20.74pt,depth=17.28pt,width=0.0pt | |
| ArtGS [liu2025artgs] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.3 | 0.2 =20.74pt,depth=17.28pt,width=0.0pt | |
| Ours | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.1 | 0.2 | 0.2 =20.74pt,depth=17.28pt,width=0.0pt | |
| Axis Pos | Ditto [Jiang_2022_ditto] | * | - | - | - | =20.74pt,depth=17.28pt,width=0.0pt | |||||||||
| PARIS [liu2023paris] | 0.9 | 1.0 | 0.0 | 0.0 | 1.4 | 3.4 | 3.4 | 1.3 | - | - | 1.1 | 0.2 | - | 0.2 | |
| PARIS* [liu2023paris] | 0.5 | 2.6 | 0.0 | 0.2 | 1.7 | 3.9 | 3.1 | 3.1 | - | - | 1.5 | 0.0 | - | 0.0 | |
| CSG-reg [weng2024digitaltwin] | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 0.0 | 0.0 | 1.5 | - | - | 0.2 | 1.1 | - | 1.1 | |
| 3Dseg-reg [weng2024digitaltwin] | - | - | 0.0 | - | - | - | - | - | - | - | - | - | - | - | |
| DigitalTwinArt [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | 0.0 | 0.0 | - | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| ArtGS [liu2025artgs] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | 0.0 | 0.0 | - | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| Ours | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | - | - | 0.0 | 0.0 | - | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| Part Motion | Ditto [Jiang_2022_ditto] | F | F | =20.74pt,depth=17.28pt,width=0.0pt | |||||||||||
| PARIS [liu2023paris] | 78.9 | 57.7 | 0.0 | 0.0 | 50.1 | 47.1 | 84.3 | 23.3 | 0.2 | 0.0 | 34.2 | 1.1 | 0.4 | 0.7 | |
| PARIS* [liu2023paris] | 75.0 | 58.5 | 0.0 | 28.8 | 64.8 | 68.1 | 67.8 | 26.5 | 0.2 | 0.0 | 39.0 | 0.0 | 0.1 | 0.1 | |
| CSG-reg [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 9.3 | 8.8 | 0.0 | 0.0 | 1.8 | 0.1 | 0.1 | 0.1 | |
| 3Dseg-reg [weng2024digitaltwin] | - | - | 0.1 | - | - | - | - | - | 0.0 | - | - | - | - | - | |
| DigitalTwinArt [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 0.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| ArtGS [liu2025artgs] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.00.0 | 0.0 | 0.0 | 0.0 | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| Ours | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| CD-s | Ditto [Jiang_2022_ditto] | =20.74pt,depth=17.28pt,width=0.0pt | |||||||||||||
| PARIS [liu2023paris] | 5.0 | 2.7 | 0.0 | 1.0 | 3.8 | 0.2 | 0.3 | 2.2 | 0.6 | 2.5 | 1.8 | 34.1 | 30.1 | 32.1 | |
| PARIS* [liu2023paris] | 5.8 | 12.0 | 0.0 | 0.3 | 13.3 | 1.9 | 0.5 | 3.7 | 0.7 | 0.8 | 3.9 | 1.5 | 2.8 | 2.1 | |
| CSG-reg [weng2024digitaltwin] | |||||||||||||||
| 3Dseg-reg [weng2024digitaltwin] | - | - | - | - | - | - | - | - | - | - | - | - | |||
| DigitalTwinArt [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.1 | 0.0 | 0.2 | 0.0 | 0.2 | 0.1 | 0.2 | 0.3 | 0.3 =20.74pt,depth=17.28pt,width=0.0pt | |
| ArtGS [liu2025artgs] | 0.3 | 0.0 | 0.0 | 0.0 | 0.3 | 0.1 | 0.2 | 0.1 | 0.0 | 0.2 | 0.1 | 0.2 | 0.3 | 0.3 =20.74pt,depth=17.28pt,width=0.0pt | |
| Ours | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.1 | 0.1 | 0.2 | 0.0 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 =20.74pt,depth=17.28pt,width=0.0pt | |
| CD-m | Ditto [Jiang_2022_ditto] | =20.74pt,depth=17.28pt,width=0.0pt | |||||||||||||
| PARIS [liu2023paris] | 7.6 | 11.2 | 0.2 | 15.2 | 40.7 | 30.7 | 5.9 | 131.7 | 9.5 | 6.1 | 25.9 | 31.7 | 382.9 | 207.3 | |
| PARIS* [liu2023paris] | 24.9 | 11.5 | 0.0 | 10.6 | 16.6 | 37.0 | 6.9 | 248.7 | 213.0 | 11.0 | 58.0 | 26.8 | 227.2 | 127.0 | |
| CSG-reg [weng2024digitaltwin] | |||||||||||||||
| 3Dseg-reg [weng2024digitaltwin] | - | - | - | - | - | - | - | - | - | - | - | - | |||
| DigitalTwinArt [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.4 | 0.2 | 0.0 | 0.1 | 0.0 | 0.1 | 0.0 | 5.0 | 2.5 =20.74pt,depth=17.28pt,width=0.0pt | |
| ArtGS [liu2025artgs] | 0.1 | 0.0 | 0.0 | 0.2 | 0.4 | 0.1 | 0.1 | 0.2 | 0.2 | 0.4 | 0.2 | 0.2 | 3.6 | 1.9 =20.74pt,depth=17.28pt,width=0.0pt | |
| Ours | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.0 | 0.0 | 0.0 =20.74pt,depth=17.28pt,width=0.0pt | |
| CD-w | Ditto [Jiang_2022_ditto] | =20.74pt,depth=17.28pt,width=0.0pt | |||||||||||||
| PARIS [liu2023paris] | 1.2 | 0.9 | 0.1 | 1.1 | 20.7 | 4.2 | 0.2 | 3.3 | 0.2 | 0.4 | 3.2 | 15.5 | 22.2 | 18.9 | |
| PARIS* [liu2023paris] | 6.4 | 4.7 | 0.0 | 0.3 | 3.6 | 5.9 | 0.2 | 2.8 | 0.1 | 0.4 | 2.4 | 1.4 | 8.2 | 4.8 | |
| CSG-reg [weng2024digitaltwin] | |||||||||||||||
| 3Dseg-reg [weng2024digitaltwin] | - | - | - | - | - | - | - | - | - | - | - | - | |||
| DigitalTwinArt [weng2024digitaltwin] | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.1 | 0.2 | 0.2 | 0.0 | 0.2 | 0.1 | 0.1 | 0.5 | 0.3 =20.74pt,depth=17.28pt,width=0.0pt | |
| ArtGS [liu2025artgs] | 0.2 | 0.0 | 0.0 | 0.0 | 0.3 | 0.0 | 0.0 | 0.1 | 0.0 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 =20.74pt,depth=17.28pt,width=0.0pt | |
| Ours | 0.0 | 0.0 | 0.0 | 0.1 | 0.0 | 0.1 | 0.0 | 0.1 | 0.0 | 0.1 | 0.0 | 0.0 | 0.1 | 0.1 =20.74pt,depth=17.28pt,width=0.0pt | |
@@
| PSNR | PARIS [liu2023paris] | 33.11 | 38.78 | 38.66 | 35.41 | 38.70 | 38.59 | 38.24 | 40.18 | 38.56 | 37.08 | 37.62 | 25.29 | 27.13 | 26.21 |
| DigitalTwinArt [weng2024digitaltwin] | 42.79 | 32.99 | 37.72 | 34.73 | 40.41 | 38.28 | 38.78 | 38.71 | 43.16 | 35.38 | 38.30 | 24.22 | 23.11 | 23.67 | |
| ArtGS [liu2025artgs] | 34.46 | 37.11 | 34.09 | 37.06 | 38.29 | 39.13 | 39.64 | 38.50 | 41.16 | 37.24 | 37.67 | 27.05 | 25.38 | 26.22 | |
@@
| SSIM | PARIS [liu2023paris] | 0.986 | 0.994 | 0.990 | 0.981 | 0.996 | 0.995 | 0.992 | 0.991 | 0.996 | 0.994 | 0.992 | 0.898 | 0.953 | 0.926 |
| DigitalTwinArt [weng2024digitaltwin] | 0.986 | 0.986 | 0.989 | 0.977 | 0.994 | 0.991 | 0.991 | 0.992 | 0.997 | 0.968 | 0.987 | 0.890 | 0.923 | 0.907 | |
| ArtGS [liu2025artgs] | 0.997 | 0.993 | 0.988 | 0.995 | 0.998 | 0.999 | 0.998 | 0.995 | 0.999 | 0.992 | 0.995 | 0.939 | 0.930 | 0.935 | |
@@
| Axis Ang 0 | Axis Ang 1 | Axis Pos 0 | Axis Pos 1 | Part Motion 0 | Part Motion 1 | CD-s | CD-m 0 | CD-m 1 | CD-w | ||
| Fridge 10489 (3 parts) | PARIS* [liu2023paris] | 34.52 | 15.91 | 3.60 | 1.63 | 86.21 | 105.86 | 8.52 | 526.19 | 160.86 | 15.00 |
| DigitalTwinArt [weng2024digitaltwin] | 0.17 | 0.09 | 0.02 | 0.00 | 0.17 | 0.11 | 0.63 | 0.44 | 0.57 | 0.89 | |
| ArtGS [liu2025artgs] | 0.02 | 0.00 | 0.00 | 0.00 | 0.02 | 0.03 | 0.62 | 0.07 | 0.18 | 0.75 | |
| Ours | 0.02 | 0.01 | 0.00 | 0.00 | 0.02 | 0.05 | 0.64 | 0.13 | 0.22 | 0.56 | |
| Storage 47254 (3 parts) | PARIS* [liu2023paris] | 43.26 | 26.18 | 10.42 | - | 79.84 | 0.64 | 8.56 | 128.62 | 266.71 | 8.66 |
| DigitalTwinArt [weng2024digitaltwin] | 0.16 | 0.80 | 0.05 | - | 0.16 | 0.00 | 0.87 | 0.23 | 0.34 | 0.99 | |
| ArtGS [liu2025artgs] | 0.01 | 0.02 | 0.01 | - | 0.01 | 0.00 | 0.78 | 0.19 | 0.27 | 0.93 | |
| Ours | 0.02 | 0.04 | 0.01 | - | 0.00 | 0.00 | 0.73 | 0.11 | 0.22 | 0.89 |
| Axis Ang | Axis Pos | Part Motion | CD-s | CD-m | CD-w | ||
| Table - 19836 (4 parts) | DigitalTwinArt [weng2024digitaltwin] | 51.74 | - | 0.80 | 10.56 | 133.87 | 0.96 |
| ArtGS [liu2025artgs] | 15.22 | - | 0.08 | 2.98 | 115.71 | 1.62 | |
| Ours | 0.01 | - | 0.00 | 0.79 | 0.27 | 1.50 | |
| Table - 25493 (4 parts) | DigitalTwinArt [weng2024digitaltwin] | 25.30 | - | 0.39 | 1.08 | 308.51 | 0.61 |
| ArtGS [liu2025artgs] | 21.59 | - | 0.05 | 0.55 | 109.13 | 0.66 | |
| Ours | 0.02 | - | 0.00 | 0.41 | 0.09 | 0.67 | |
| Storage - 45503 (4 parts) | DigitalTwinArt [weng2024digitaltwin] | 41.77 | 3.18 | 26.60 | 2.40 | 176.78 | 0.98 |
| ArtGS [liu2025artgs] | 4.86 | 2.82 | 10.28 | 1.29 | 3.20 | 4.86 | |
| Ours | 0.02 | 0.01 | 0.09 | 0.70 | 0.10 | 0.97 | |
| Storage - 41083 (5 parts) | DigitalTwinArt [weng2024digitaltwin] | 61.11 | 5.30 | 34.48 | 2.39 | 355.04 | 1.16 |
| ArtGS [liu2025artgs] | 1.08 | 1.65 | 10.28 | 1.98 | 120.15 | 2.05 | |
| Ours | 0.03 | 0.01 | 0.06 | 0.94 | 0.16 | 6.77 | |
| Storage - 46145 (7 parts) | DigitalTwinArt [weng2024digitaltwin] | 64.77 | 4.30 | 25.68 | 1.58 | 400.71 | 0.84 |
| ArtGS [liu2025artgs] | 8.49 | 3.30 | 12.29 | 1.54 | 359.14 | 2.30 | |
| Ours | 0.07 | 0.00 | 0.06 | 0.52 | 0.09 | 1.58 |